Which number is divisible by 2 and by 3?

(A) 112 (B) 4308 (C) 6122 (D) 23,451 (E) 701,456

Answer by mananth(12270)

any even number is divisible by 2.
When you add the digits of any number the total if divisible by 3 then the whole number s divisible by 3.
For example 2646 2+6+4+6 = 18 again add these two digits . 8+1 =9 . 9 is divisible by 3.
So 2646 is divisible by 3.
In your problem you have to select an even number and then test for divisiblity of 3.
.
.
4308 is an even number. 4+3 +0+8 = 15. --- 5+1 = 6. 6 is divisible by 3 . so 4308 is divisible by 3 also.